OpenStudy (anonymous):

The rational function has a y-intercept of What is the equation for this function?

OpenStudy (anonymous):

http://mathworld.wolfram.com/RationalFunction.html

OpenStudy (anonymous):

I'm not really getting it :/ can u explain it please

OpenStudy (anonymous):

horizontal asymptote is \(x=-3\) so the denominator is probably \(x+3\)

OpenStudy (anonymous):

vertical asymptote is at \(y=5\) so it is something like \[y=\frac{5x}{x+3}\] but since it crosses the \(y\) axis at \((0,7)\) you have something like \[y=\frac{5x+7}{x+3}\]

OpenStudy (anonymous):

no that was wrong!!

OpenStudy (anonymous):

sorry. since \(0\) gets you 7 you can solve \[\frac{5x+b}{x+3}=7\]

OpenStudy (anonymous):

ohh ok so this is the final answer o.O sorry a little confused

OpenStudy (anonymous):

jeez i am an idiot, sorry let me try again

OpenStudy (anonymous):

we know it looks like \[y=\frac{5x+b}{x+3}\] because of the asymptotes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we know that if \(x=0\) we get \(y=7\) so solve \[\frac{b}{3}=7\] and get \(b=21\)

OpenStudy (anonymous):

therefore it should be \[y=\frac{5x+21}{x+3}\]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

damn i mean \[f(x)=\frac{5x+21}{x+3}\] http://www.wolframalpha.com/input/?i=%285x%2B21%29%2F%28x%2B3%29+domain+-9..9+

OpenStudy (anonymous):

nice head on ;D thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thanks :D <3